OpenStudy (anonymous):

Algebra I Help!! 12.) a^2 + 5a + 4 over a^3 divided by a^2 + 3a + 2 over a^2 - 2a. 14.) (12x^2 + 9x - 7) divided by 3x.

OpenStudy (anonymous):

It is Multiplying & Dividing Rational Expressions & Dividing Polynomials... So please help, i am behinnd and need help! Thank You:)

OpenStudy (anonymous):

okay... so what does over a^3 mean?

OpenStudy (anonymous):

a^2 + 5a + 4/a^3 divided by a^2 + 3a + 2/a^2 - 2a

OpenStudy (espex):

Is your question this? \[\frac{\frac{a^2 + 5a + 4}{a^{3}}}{\frac{a^2 + 3a + 2}{a^2 - 2a}}\]

OpenStudy (anonymous):

Yes

OpenStudy (espex):

Any time you have a fraction, divided by a fraction, you can rewrite it as the numerator fraction multiplied by the inverse of the denominator fraction. like this: \[\frac{\frac{1}{2}}{\frac{2}{4}}\rightarrow\frac{1}{2}\times\frac{4}{2}\]

OpenStudy (espex):

So you can rewrite your problem like this: \[\frac{a^2 + 5a + 4}{a^{3}}\times\frac{a^2 - 2a}{a^2 + 3a + 2}\]

OpenStudy (espex):

Now if you factor each of these you can simplify before you multiply them.